# Machine-assisted classification of potential biosignatures in earth-like exoplanets using low signal-to-noise ratio transmission spectra

### v1

@@ Abstract
The search for atmospheric biosignatures in earth-like exoplanets is one of the most pressing challenges in observational astrobiology. Detecting biogenic gases in terrestrial planets requires high resolution and long integration times. In this work, we developed and tested a machine-learning general methodology, intended to classify transmission spectra with low Signal-to-Noise Ratio according to their potential to contain biosignatures. For that purpose, we trained a set of models capable of classifying noisy transmission spectra, as having methane, ozone, and/or water (multilabel classification), or simply as being interesting for follow-up observations (binary classification). The models were trained with $\sim 10^{6}$ synthetic spectra of planets similar to TRAPPIST-1 e which were generated with the package MultiREx especially developed for this work. The trained algorithms correctly classified test planets with transmission spectra having SNR<6 and containing methane and/or ozone at mixing ratios similar to those of modern and Proterozoic Earth. Tests on realistic synthetic spectra based on the current Earth’s atmosphere show at least one of our models would classify as likely having biosignatures and using only one transit, most of the inhabited terrestrial planets observed with the JWST/NIRSpec PRISM around M-dwarfs located at distances similar or smaller than that of TRAPPIST-1 e. The implication of this result for the designing of observing programs and future surveys is enormous since machine-assisted strategies similar to those presented here could significantly optimize the usage of JWST resources for biosignature searching, while maximizing the chances of a real discovery after dedicated follow-up observations of promising candidates.
@@
pubyear: 2024pagerange: Machine-assisted classification of potential biosignatures in earth-like exoplanets using low signal-to-noise ratio transmission spectraMachine-assisted classification of potential biosignatures in earth-like exoplanets using low signal-to-noise ratio transmission spectra
@@ 1 Introduction
One of the strategies conceived to search for life beyond Earth involves detecting signals on exoplanets that may be closely associated with life and have a low probability of being abiotic in origin. We call these signals, biosignatures (see e.g. Schwieterman et al. 2018; Des Marais et al. 2008 and references therein). Currently, our instrumental capabilities allow us to detect biosignatures that have a global impact, especially those concentrated in the atmosphere and observable through spectroscopy (see e.g. Schwieterman et al. 2018).
@@
However, detecting biosignatures on rocky exoplanets within the habitable zone of their stars poses significant challenges due to the low signal-to-noise ratio resulting from the smaller relative radii of the star-to-planet. M-dwarfs offer an opportunity to study the atmospheres of Earth-like planets. Although these stars may exhibit high X-ray and ultraviolet (XUV) activity in their early stages, secondary atmospheres may remain stable as they age (France et al., 2020). The fact is that M-dwarfs allow for greater transit depth compared to other types of stars (Wunderlich et al., 2019) increasing the chances of detecting and analysing exoplanetary atmospheres. Instruments such as JWST demonstrate our ability to reliably detect atmospheric species on exoplanets around M-dwarfs, such as possible H2-rich atmospheres (Madhusudhan et al., 2023; Benneke et al., 2024), and even showcase the capability to identify, as the most likely scenario, Earth-like atmospheres such as N2- or CO2-dominated atmospheres (Cadieux et al., 2024).
@@
Recent simulations on the detectability of an Earth-like atmosphere using the James Webb Space Telescope (JWST) (Barstow & Irwin, 2016; Wunderlich et al., 2019; Lin et al., 2021; Lustig-Yaeger et al., 2023) have revealed a challenging scenario. To detect robust biosignatures such as ozone ($O_{3}$), requires a large number of transits (e.g., up to 200 transits in the case of TRAPPIST-1 e; see Lin et al. 2021) to achieve statistically significant detections. Despite this challenge, detecting methane ($CH_{4}$) and water vapour ($H_{2}O$) presents a promising opportunity. Studies have demonstrated that using a reasonable number of transits, the presence of these atmospheric species, which are typically associated with a global biosphere, can be retrieved (Wunderlich et al., 2019; Lin et al., 2021; Lustig-Yaeger et al., 2023). However, it is important to note that $CH_{4}$ is a less robust biosignature compared to $O_{3}$ (for more details on the robustness of ozone as a biosignature, see subsection 4.2).
@@
Given the almost prohibitive cost of performing such an observational campaign, a more relevant strategy may be allocating JWST time to conduct a low signal-to-noise ratio (SNR) survey. Although this may not allow for statistically significant retrievals, it would at least enable planning for future follow-up observations of interesting targets with current and future more powerful telescopes (e.g., ELT, LUVOIR, HabEx, Roman, ARIEL).
@@
In this paper, we propose and numerically test the methods and tools to support such an strategy. For this purpose we design and test a set of machine learning (ML) methods and tools aimed at labelling low SNR spectra that have the potential to contain interesting biosignatures. It is important to stress that the tools presented here are not designed to perform a retrieval of the abundances of chemical species, but to identify interesting candidates for follow-up observations.
@@
The use of Machine Learning (ML) for analysing exoplanet spectra is gaining increasing interest. Multiple algorithms have been developed and trained on diverse datasets and objectives (Marquez-Neila et al., 2018; Soboczenski et al., 2018; Zingales & Waldmann, 2018; Cobb et al., 2019; Nixon & Madhusudhan, 2020). Primarily, these algorithms have been used to replace the retrieval process based on Bayesian methods, which is characterised by its high computational cost (Munsaket et al., 2021; Ardévol Martínez et al., 2022; Vasist et al., 2023; Ardévol Martínez et al., 2024). In Table 1 we summarise some of the strategies that have been devised for performing atmospheric composition retrieval and are available in the literature.
@@
Forestano et al. (2023) LOF, 1CSVM (Unsupervised ML, Outlier detection) Hot Jupiters
@@
Hayes et al. (2020) k-means, PCA (Unsupervised classification) Jupiter-like planets
@@
Additionally, the capacity of ML to support parts of the retrieval process, such as suggesting priors (Hayes et al., 2020), identifying molecules (Waldmann, 2016), or using neural network parametrization of pressure-temperature profiles for better efficiency and physical consistency (Gebhard et al., 2024), or even replacing the spectrum generation process through radiative transfer, has been explored (Himes et al., 2022). Conversely, unsupervised learning techniques have also been proposed for analysing spectroscopic data and detecting anomalous chemical compositions (Forestano et al., 2023; Matchev et al., 2022a). ML has been used to assess molecule detection capabilities in various instruments of the James Webb Space Telescope (JWST), demonstrating how machine learning techniques have become robust alternatives to classical retrieval methods (Guzmán-Mesa et al., 2020).
@@
Furthermore, these studies have primarily focused on gaseous planets, especially Hot Jupiters like WASP-12. There is a clear minority of studies on rocky and Earth-like planets, indicating a significant gap in the training data available for these types of planets. Nonetheless, recent research has started to consider brown dwarfs, as seen in the work of Ardévol Martínez et al. (2024) and Lueber et al. (2023). The implemented regression techniques commonly include Random Forest and various types of neural networks. Additionally, unsupervised learning techniques are particularly focused on clustering and anomaly detection.
@@
As pointed out before, our proposal is not to perform the full retrieval, which, even using ML techniques, requires high SNR signals. This is demonstrated by the fact that many of the previously cited works are related to giant planets. But, if we can take low SNR spectra and label interesting Earth-like planets, we can focus our search on the best candidates.
This paper is organized as follows. In section 2, we describe the ML methods used in this paper, particularly introducing the language and quantities used in classification-supervised machine learning. To test our methods, we focus on a very interesting target, namely TRAPPIST-1 e, which is described in section 3. The section 4 is dedicated to describing the most interesting species on which our numerical experiments are focused. The methods, tools, and preliminary results are detailed in section 5 and section 6. Contextualization of the results based on the required transits using the NIRSpec MIRI instrument of JWST is discussed in section 7. Our algorithms are tested using realistic spectra of Earth in section 8. The limitations of our approach and the training data are discussed in section 9. Finally, the conclusions of our numerical experiments are drawn in section 10. All the tools and data developed for this paper are publicly available, with access information provided in section 11.
@@ 2 Supervised Machine Learning for biosignatures detection
Machine learning for classification (MLC) is an approach to building a predictive model that assigns a label or category to a data point or instance based on a set of known features (see e.g., Géron 2023). For example, we aim to label planets based on the presence of biosignatures using simple categories: $O_{3}/\text{\tt No-}O_{3}$, $CH_{4}/\text{\tt No-}CH_{4}$.
In MLC, a training set consists of labelled observations (e.g., a list of planets with and without biosignatures), where each observation comprises features and a classification label. The simplest form is binary classification, where an instance either belongs to a category or does not. The goal is for the classificator model to accurately predict the classification label for new, unlabelled data, enabling decisions based on the classification of such data. Our objective is for the trained algorithm to identify which planets have a specific biosignature and which do not, using a low SNR spectrum.
@@ 2.1 Random forest for biosignature classification
Many algorithms have been devised for MLC. Here, we focus on Random Forest (RF), a family of methods that is not only simple enough but also particularly well-suited to our requirements. RF demonstrates robustness to noise by maintaining high accuracy through the aggregation of predictions from multiple trees, thereby reducing the impact of noisy instances. Additionally, RF is inherently resistant to overfitting due to the Law of Large Numbers, which ensures the convergence of the model’s error as more trees are added (Breiman, 2001). This characteristic makes RF especially effective for complex datasets with numerous variables, each contributing marginally to the prediction, providing a robust solution for high-dimensional data classification.
In our case, RF takes signals from various spectral bands (these are the features) and creates decision trees based on the values of these specific signals. Each decision tree in the forest can use multiple features simultaneously to make decisions, determining if these signals contribute to identifying a biosignature.
@@ 2.2 Metrics for classification
In classification problems, the performance of a machine learning algorithm is not only measured in terms of how many planets were correctly classified (“True” cases), i.e. the accuracy or how many were not (“False” cases), i.e. the errors. It is also required to categorize the classification results into positives and negatives, either True Positives (TP) or True Negatives (TN), which are related to accuracy; and into False Positives (FP) or False Negatives (FN), which are the error. This categorization is summarised in what is called a confusion matrix (Kelleher et al., 2015).
In Figure 1 we schematically represent the categorization in a confusion matrix and, more importantly, illustrate what each case means in the context of potentially biosignature-bearing exoplanets. In the diagram, we introduce the category interesting to distinguish planets that deserve follow-up observations or in-depth analysis. We should recall again that is the focus of this work: we do not aim at detecting biosignatures using ML but at labeling planets that are interesting or not.
Figure 1: Schematic representation of the confusion matrix used in ML classification but for the context of biosignature searching. Instances (planets) in the first raw have biosignatures, while those in the second one are devoid of them. An algorithm labels instances (planets) in the first column as potentially having biosignatures (interesting), while the algorithm does not consider those in the second column as interesting.
Different types of errors carry specific implications for biosignature searching. For example, mistakenly classifying a planet that contains biosignatures as not interesting (FN), could mean the loss of valuable research opportunities. In contrast, erroneously considering a planet without biosignatures as interesting (FP) could lead to inefficient resource allocation. Therefore, understanding the difference between these types of errors is essential for evaluating the accuracy of a model and for adjusting decision and exploration strategies based on the model’s classification outcomes (see section 6).
Given the limitations of using accuracy alone, i.e. number of right predictions divided by the total number of predictions, especially when different errors have disproportionately large impacts, more nuanced metrics become indispensable (Kelleher et al., 2015). For instance, recall which is defined as
@@
measures the model’s efficiency in detecting all relevant instances. Ideally, a $\text{Recall}=1$ means that the algorithm correctly classifies all planets actually having biosignatures (first row in Figure 1) as interesting exoplanets. In this case, the follow-up observations will be maximally successful. For this reason, we call Recall a wasting metric: when Recall decreases, we are excluding promising planets from future observational campaigns and wasting research opportunities.
@@
that assesses the accuracy of interesting predictions. This metric is more difficult to interpret since it focuses on the classification rather than the actual presence of biosignatures. When the number of planets without biosignatures classified as interesting (FP) decreases, the Precision increases. Therefore, we call Precision a time-saving metric, as it helps minimize the time spent studying planets without biosignatures.
In summary, we expect that an effective algorithm applied to a given dataset would have a high value for both metrics, i.e., it will be minimally wasteful and maximally time-saving.
But there is a third metric, one quantifying another aspect of the problem: the discovery opportunity. For this purpose, we use the F1 Score, defined as:
@@
All of the previous metric are focused on positive classifications. However, for the specific purpose of searching for biosignatures, and especially for studying the potential confusion that can arise when an algorithm is designed to detect a specific biosignature in the presence of other molecules, we should use metrics focused on the negative cases. Thus, for example, the True Negative Rate (TNR):
@@
We call this metric a confusion metric, in the sense that when the TNR decreases, the algorithm labels planets without biosignatures as interesting. This confusion can arise, for instance, when a biologically irrelevant molecule is “identified” by the algorithm as a biosignature. Essentially, TNR is similar to recall but focuses on the negative labels, ensuring that planets without biosignatures are correctly identified as not interesting. Therefore, TNR can also be understood as a negative recall.
Since understanding the definition, meaning and application of a given metric could be very confusing, we have summarized them in Table 2 with special attention on highlight their role in context of biosignature search.
Table 2: Summary of metrics used in the evaluation of biosignature detection algorithms. The table includes the metric name, a nickname describing its utility, its specific use in the context of evaluating planets for biosignatures, and extreme cases illustrating the possible values and interpretations of each metric.
@@
Recall Wasting metric When small we are excluding promising planets from future observational campaigns, and wasting research opportunities Recall = 1: all planets with biosignatures are labelled as interesting Recall = 0: no planets with biosignatures are labelled as interesting
Precision Time-saving metric Helps to minimize the time spent studying planets without biosignatures Precision = 1: All planets labelled as interesting truly have biosignatures. Precision = 0: Any planet is labelled as interesting even without biosignatures. Note: Even if the false positives are at their maximum, meaning all non-interesting planets are incorrectly labelled as interesting, the precision is not zero as long as there are true positives. This occurs because precision is calculated as the ratio of true positives to the sum of true positives and false positives. Thus, the presence of any correctly identified interesting planets ensures that precision remains above zero.
@@
TNR Confusion metric Helps to ensure that planets without biosignatures are not mislabelled as interesting, thereby avoiding wasted resources on unpromising candidates TNR = 1: all non-interesting planets are correctly identified as not interesting TNR = 0: all non-interesting planets are incorrectly labelled as interesting
Hamming Loss Error rate metric It measures the fraction of incorrectly predicted molecules, indicating the algorithm’s deficiency in misclassifying or omitting the presence of molecules in general. Hamming Loss = 0: All predicted labels (biosignatures) are correct. Hamming Loss = 1: All predicted labels (biosignatures) are incorrect.
Exact Match Ratio Perfect match metric Measures how well the algorithm classifies all biosignatures for planets correctly at the same time. This metric is very demanding as it requires perfect classification for all labels. Higher values indicate better performance. Exact Match Ratio = 1: The algorithm correctly assigns all biosignatures to all planets. Exact Match Ratio = 0: The algorithm fails to assign the correct set of biosignatures to any planet.

### v4

@@ Abstract
The search for atmospheric biosignatures in Earth-like exoplanets is one of the most pressing challenges in observational astrobiology. Detecting biogenic gases in terrestrial planets requires high-resolution observations and long integration times. In this work, we developed and tested a general machine-learning methodology designed to classify transmission spectra with low Signal-to-Noise Ratio (SNR) according to their potential to contain biosignatures or bioindicators. To achieve this, we trained a set of models capable of classifying noisy transmission spectra (including stellar contamination) as containing methane, ozone, and/or water (multilabel classification), or simply as being interesting for follow-up observations (binary classification). The models were trained using $\sim 10^{7}$ synthetic spectra of planets similar to TRAPPIST-1 e, generated with the package MultiREx. The trained algorithms correctly classified most of the test planets with transmission spectra having an SNR as low as 4, containing methane and/or ozone at mixing ratios similar to those of modern and Proterozoic Earth. Tests on realistic synthetic spectra, based on the current Earth’s atmosphere, indicate that some of our models would classify most inhabited terrestrial planets observed with JWST/NIRSpec PRISM around M-dwarfs at distances similar to or smaller than that of TRAPPIST-1 e as likely to contain bioindicators, using 4–10 transits. These results have significant implications for the design of observing programs and future campaigns. Machine-assisted strategies, such as the one presented here, could greatly optimize the use of JWST resources for biosignature and bioindicator searches, while maximizing the chances of a real discovery through dedicated follow-up observations of promising candidates.
@@
pubyear: 2025pagerange: Machine-assisted classification of potential biosignatures in Earth-like exoplanets using low signal-to-noise ratio transmission spectraMachine-assisted classification of potential biosignatures in Earth-like exoplanets using low signal-to-noise ratio transmission spectra
@@ 1 Introduction
One of the strategies conceived to search for life beyond Earth involves detecting signals on exoplanets that may be closely associated with life and have a low probability of being abiotic in origin. We call these signals, biosignatures (see e.g. Schwieterman et al. 2018; Des Marais et al. 2008 and references therein). Currently, our instrumental capabilities allow us to detect biosignatures that have a global impact, especially those concentrated in the atmosphere and observable through spectroscopy (see e.g. Schwieterman et al. 2018). Furthermore, in cases where the likelihood of an abiotic origin is not negligible, the concept of a bioindicator can be utilised to distinguish molecular species that are more likely to be produced through biological processes. Bioindicators, while not definitive evidence of life on their own, provide valuable complementary clues that help to discern biological activity from abiotic processes.
@@
However, detecting biosignatures on rocky exoplanets within the habitable zone of their stars poses significant challenges due to the low signal-to-noise ratio resulting from the smaller relative radii of the star-to-planet. M-dwarfs offer an opportunity to study the atmospheres of Earth-like planets. Although these stars may exhibit high X-ray and ultraviolet (XUV) activity in their early stages, secondary atmospheres may remain stable as they age (France et al., 2020). It is well established that M-dwarfs exhibit greater transit depths compared to other stellar types (Wunderlich et al., 2019) increasing the chances of detecting and analysing exoplanetary atmospheres. Instruments such as the James Webb Space Telescope (JWST) demonstrate our ability to reliably detect atmospheric species on exoplanets around M-dwarfs, such as possible H2-rich atmospheres (Madhusudhan et al., 2023; Benneke et al., 2024), and even showcase the capability to identify, as the most likely scenario, Earth-like atmospheres such as N2- or CO2-dominated atmospheres (Cadieux et al., 2024).
@@
Recent simulations on the detectability of an Earth-like atmosphere using the JWST (Barstow & Irwin, 2016; Wunderlich et al., 2019; Lin et al., 2021; Lustig-Yaeger et al., 2023) have revealed a challenging scenario. To detect robust bioindicators such as ozone (O3) (for the distinction between bioindicators and biosignatures, see section 4), a large number of transits is required (e.g., up to 200 transits in the case of TRAPPIST-1 e; see Lin et al. 2021) to achieve statistically significant detections. Despite this challenge, detecting methane (CH4) and water vapour (H2O) presents a promising opportunity. Studies have demonstrated that, using a reasonable number of transits, the presence of these atmospheric species, which are typically associated with a global biosphere, can be retrieved (Wunderlich et al., 2019; Lin et al., 2021; Lustig-Yaeger et al., 2023). However, it is important to note that CH4 is a less robust bioindicator compared to O3 (for more details on the robustness of ozone as a bioindicator, see subsection 4.2).
@@
Considering the substantial resource requirements associated with conducting such observational campaigns, a more effective strategy may be to allocate JWST time to conduct a low signal-to-noise ratio (SNR) survey. Although this may not allow for statistically significant retrievals, it would facilitate planning future follow-up observations of promising candidate targets with current and future, more powerful telescopes (e.g., ELT, LUVOIR, HabEx, Roman, ARIEL).
@@
In this paper, we propose and numerically test methods and tools to support such a strategy. For this purpose, we design and test a set of machine learning (ML) methods and tools aimed at labelling low SNR spectra that may contain interesting biosignatures or bioindicators. It is important to stress that the tools presented here are not designed to perform a retrieval of the abundances of chemical species, but to identify interesting candidates for follow-up observations.
@@
The use of ML in analysing exoplanet spectra has garnered significant attention in recent literature. Multiple algorithms have been developed and trained on diverse datasets and objectives (Marquez-Neila et al., 2018; Soboczenski et al., 2018; Zingales & Waldmann, 2018; Cobb et al., 2019; Nixon & Madhusudhan, 2020). Primarily, these algorithms have been employed as alternatives to the computationally intensive Bayesian retrieval methods (Munsaket et al., 2021; Ardévol Martínez et al., 2022; Vasist et al., 2023; Ardévol Martínez et al., 2024). In Table 1 we summarise some of the strategies that have been devised for performing atmospheric composition retrieval and are available in the literature.
@@
Forestano et al. (2023) Local Outlier Factor, One-Class Support Vector Machine (Unsupervised Machine Learning, Outlier Detection) Hot Jupiters
@@
Hayes et al. (2020) K-means, Principal Component Analysis (Unsupervised Classification) Jupiter-like planets
@@
Additionally, the capacity of ML to support parts of the retrieval process – such as suggesting priors (Hayes et al., 2020), identifying molecules (Waldmann, 2016), using neural network parametrisation of pressure-temperature profiles for better efficiency and physical consistency (Gebhard et al., 2024), or even replacing the spectrum generation process through radiative transfer – has been explored (Himes et al., 2022). Conversely, unsupervised learning techniques have also been proposed for analysing spectroscopic data and detecting anomalous chemical compositions (Forestano et al., 2023; Matchev et al., 2022a). ML has been used to assess molecule detection capabilities in various instruments of the JWST, demonstrating how machine learning techniques have become robust alternatives to classical retrieval methods (Guzmán-Mesa et al., 2020).
@@
Furthermore, these studies have primarily focused on gaseous planets, especially Hot Jupiters like WASP-12. Only a minority of studies focus on rocky and Earth-like planets, indicating a significant gap in the training data available for these types of planets. Nonetheless, recent research has started to consider brown dwarfs, as seen in the work of Ardévol Martínez et al. (2024) and Lueber et al. (2023). The implemented regression techniques commonly include Random Forest and various types of neural networks. Additionally, unsupervised learning techniques are particularly focused on clustering and anomaly detection.
@@
As pointed out before, our proposal is not to perform the full retrieval, which—even using ML techniques—requires high SNR signals. This is demonstrated by the fact that many of the previously cited works are related to giant planets. However, if we can obtain low SNR spectra and identify intriguing Earth-like planets, we can concentrate our search on the most promising candidates.
This paper is organised as follows. In section 2 we describe the ML methods used in this paper, particularly introducing the language and quantities used in classification-supervised machine learning. To test our methods, we focus on a very interesting target, namely TRAPPIST-1 e, which is described in section 3. section 4 is dedicated to describing the most interesting species on which our numerical experiments are focused. Furthermore, section 5 provide a comprehensive discussion of stellar contamination—an effect that must be taken into account when assessing how it affects the observed transit spectrum by introducing spurious features that might be mistakenly attributed to the planet—and outline the autoencoder approach we use to address it. The methods, tools, and preliminary results are detailed in section 6 and section 7. Contextualisation of the results based on the required transits using the NIRSpec MIRI instrument of JWST is discussed in section 8. Our algorithms are tested using realistic spectra of Earth in section 9. The limitations of our approach and the training data are discussed in section 10. Finally, the conclusions of our numerical experiments are drawn in section 11. All the tools and data developed for this paper are publicly available, with access information provided in Data availability and reproducibility.
@@ 2 Supervised Machine Learning for biosignatures detection
[…]
In MLC, a training set consists of labelled observations (e.g., a list of planets with and without biosignatures/bioindicators), where each observation comprises features and a classification label. The simplest form is binary classification, where an instance either belongs to a category or does not. The goal is for the classifier model to accurately predict the classification label for new, unlabelled data, enabling decisions based on the classification of such data. Our objective is for the trained algorithm to identify which planets have a specific biosignature and which do not, using a low SNR spectrum.
@@ 2.1 Random forest for biosignature classification
Many algorithms have been devised for MLC. Here, we focus on Random Forest (RF), a family of methods that is not only simple enough but also particularly well-suited to our requirements. RF demonstrates robustness to noise by maintaining high accuracy through the aggregation of predictions from multiple trees, thereby reducing the impact of noisy instances. Additionally, RF is inherently resistant to overfitting due to the Law of Large Numbers, which ensures the convergence of the model’s error as more trees are added (Breiman, 2001). This characteristic makes RF especially effective for complex datasets with numerous variables, each contributing marginally to the prediction, thus providing a robust solution for high-dimensional data classification.
In our case, RF takes signals from various spectral bands (these are the features) and creates decision trees based on the values of these specific signals. Each decision tree in the forest can use multiple features simultaneously to make decisions, determining whether these signals contribute to identifying a biosignature.
@@ 2.2 Metrics for classification
In classification problems, the performance of a machine learning algorithm is not only measured in terms of how many planets were correctly classified (“True” cases), i.e. the accuracy, or how many were not (“False” cases), i.e. the errors. It is also necessary to categorise the classification results into positives and negatives: either True Positives (TP) or True Negatives (TN), which are related to accuracy; and into False Positives (FP) or False Negatives (FN), which correspond to errors. This categorization is summarised in what is called a confusion matrix (Kelleher et al., 2015).
In Figure 1 we schematically represent the categorization in a confusion matrix and, more importantly, illustrate what each case means in the context of potentially biosignature-bearing exoplanets. In the diagram, we introduce the category interesting to distinguish planets that deserve follow-up observations or in-depth analysis.We should recall that this is the focus of this work: we do not aim to detect biosignatures/bioindicators using ML but rather to label planets as interesting or not.
Figure 1: Schematic representation of the confusion matrix employed for ML classification in the context of biosignature searching. Instances (planets) in the first row possess biosignatures/bioindicators, while those in the second row do not. An algorithm labels instances (planets) in the first column as potentially having biosignatures (i.e., interesting), whereas those in the second column are not considered interesting. This framework illustrates the different classification outcomes, including true positives (TP), false positives (FP), true negatives (TN), and false negatives (FN), in the search for biosignatures/bioindicators.
Different types of errors carry specific implications for biosignature searching. For example, mistakenly classifying a planet that contains biosignatures/bioindicators as not interesting (FN), could mean the loss of valuable research opportunities. In contrast, erroneously considering a planet without biosignatures/bioindicators as interesting (FP) could lead to inefficient resource allocation. Therefore, understanding the difference between these types of errors is essential for evaluating the accuracy of a model and for adjusting decision and exploration strategies based on the model’s classification outcomes (see section 7).
Given the limitations of using accuracy alone, i.e. number of correct predictions divided by the total number of predictions, especially when different errors have disproportionately large impacts, more nuanced metrics become indispensable (Kelleher et al., 2015). For instance, recall which is defined as
@@
measures the model’s efficiency in detecting all relevant instances. Ideally, a $\text{Recall}=1$ means that the algorithm correctly classifies all planets that have biosignatures/bioindicators (first row in Figure 1) as interesting exoplanets. In this case, the follow-up observations will be maximally successful. For this reason, we call Recall a wasting metric: when Recall decreases, we are excluding promising planets from future observational campaigns and wasting research opportunities.
@@
that assesses the accuracy of interesting predictions. This metric is more difficult to interpret since it focuses on the classification rather than the actual presence of biosignatures/bioindicators. When the number of planets without biosignatures/bioindicators classified as interesting (FP) decreases, Precision increases. Therefore, we call Precision a time-saving metric, as it helps minimise the time spent studying planets without biosignatures/bioindicators.
In summary, we expect that an effective algorithm applied to a given dataset would have high values for both metrics, i.e., it will be minimally wasteful and maximally time-saving.
But there is a third metric, one that quantifies another aspect of the problem: the discovery opportunity. For this purpose, we use the F1 Score, defined as:
@@
All of the previous metrics are focused on positive classifications. However, for the specific purpose of searching for biosignatures/bioindicators, and especially for studying the potential confusion that can arise when an algorithm is designed to detect a specific biosignature in the presence of other molecules, we should use metrics focused on the negative cases. Thus, for example, the True Negative Rate (TNR):
@@
We call this metric a confusion metric, in the sense that when the TNR decreases, the algorithm labels planets without biosignatures/bioindicators as interesting. This confusion can arise, for instance, when a biologically irrelevant molecule is “identified” by the algorithm as a biosignature. Essentially, TNR is similar to recall but focuses on the negative labels, ensuring that planets without biosignatures/bioindicators are correctly identified as not interesting. Therefore, TNR can also be understood as a negative recall.
Since understanding the definition, meaning and application of a given metric can be very confusing, we have summarised them in Table 2 with special attention on highlighting their role in the context of biosignature search.
Table 2: Summary of metrics used in the evaluation of biosignature detection algorithms. The table includes the metric name, a nickname describing its utility, its specific use in the context of evaluating planets for biosignatures/bioindicators, and extreme cases illustrating the possible values and interpretations of each metric.
@@
Recall Wasting metric When small we are excluding promising planets from future observational campaigns, and wasting research opportunities Recall = 1: all planets with biosignatures/bioindicators are labelled as interesting Recall = 0: no planets with biosignatures/bioindicators are labelled as interesting
Precision Time-saving metric Helps to minimize the time spent studying planets without biosignatures/bioindicators Precision = 1: All planets labelled as interesting truly have biosignatures/bioindicators. Precision = 0: Any planet is labelled as interesting even without biosignatures/bioindicators. Note: Even if the false positives are at their maximum, meaning all non-interesting planets are incorrectly labelled as interesting, the precision is not zero as long as there are true positives. This occurs because precision is calculated as the ratio of true positives to the sum of true positives and false positives. Thus, the presence of any correctly identified interesting planets ensures that precision remains above zero.
@@
TNR Confusion metric Helps to ensure that planets without biosignatures/bioindicators are not mislabelled as interesting, thereby avoiding wasted resources on unpromising candidates TNR = 1: all non-interesting planets are correctly identified as not interesting TNR = 0: all non-interesting planets are incorrectly labelled as interesting
Hamming Loss Error rate metric It measures the fraction of incorrectly predicted molecules, indicating the algorithm’s deficiency in misclassifying or omitting the presence of molecules in general. Hamming Loss = 0: All predicted labels (biosignatures/bioindicators) are correct. Hamming Loss = 1: All predicted labels (biosignatures/bioindicators) are incorrect.
Exact Match Ratio Perfect match metric Measures how well the algorithm classifies all biosignatures/bioindicators for planets correctly at the same time. This metric is very demanding as it requires perfect classification for all labels. Higher values indicate better performance. Exact Match Ratio = 1: The algorithm correctly assigns all biosignatures/bioindicators to all planets. Exact Match Ratio = 0: The algorithm fails to assign the correct set of biosignatures/bioindicators to any planet.